OpenStudy (anonymous):

-v+5+6v=1+5v+3 it is no solution but i need to know this steps to get that answer please help

OpenStudy (anonymous):

Do you know how to move stuff around by doing the opposite operation to it?

OpenStudy (anonymous):

no i dont

OpenStudy (anonymous):

Okay your goal is to isolate the variable. Lets start with a simple example. If you have 5+x=10 and you want to figure out what x is you just take away 5 from each side and you get get x=5 understand? So start by doing the opposite operations to get the numbers on one side and the numbers with variables next to them on the other side. It doesn't matter which side you choose for each. Try this first and I will help you on the last step.

OpenStudy (anonymous):

sorry it sounds sinple but then i look at my problem and my brain goes goodnight

OpenStudy (anonymous):

The lines are parallel?

OpenStudy (anonymous):

It's okay just relax and take it slow you'll get there. huh what lines timo? Am I confused?

OpenStudy (anonymous):

For questions like these, just take it step by step and with practice, it will be easier. -v+5+6v=1+5v+3 Before moving the numbers left and right of the equation, try and simplify it on their own first. If it helps, you can rearrange it also to become the following, 6v-v +5 = 5v +1 +3 Now you can solve a bit and remove some numbers, making lesser terms. 5v +5 = 5v +4 Now you notice that you have 5v on both sides, so they cancel out each other, leaving 5=4. Now we know that 5 is not equals to 4. So we can immediately conclude that there is no value of v that will me this equation correct. Thus no solution for v. Hope i made it easier for you to understand :D

OpenStudy (anonymous):

Okay good I thought this was just algebra haha.

OpenStudy (anonymous):

there is no solution cuzz if you were to graph -v+5+6v 1+5v+3 they would be parallel.

OpenStudy (anonymous):

Just a side note :P How to express it graphically

OpenStudy (anonymous):

thats crazy that you immediately imagined it visually

OpenStudy (anonymous):

yes that does makes some sence to me tiaph and its regastering thank you

OpenStudy (anonymous):

When I hear no solution I hear no intersection :P